OpenStudy (kendricklamar2014):

You have five $1 bills, four $5 bills, six $10 bills, and three $20 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill, What is P($1, then $10)? A. 11/35 B. 5/51 C. 5/54 D. 193/306

OpenStudy (kendricklamar2014):

@kropot72

OpenStudy (kendricklamar2014):

@Maria195

OpenStudy (anonymous):

Don't worry mah friend. Let me check in my previous lessons.

OpenStudy (anonymous):

Aww my numbers are different than yours X(

OpenStudy (kendricklamar2014):

:(

OpenStudy (kropot72):

The probability of choosing a $1 bill first is 5/18. the probability of choosing a $10 bill next is 6/17. The required probability is: \[P($1\ then\ $10)=\frac{5\times6}{18\times17}=you\ can\ calculate\]

OpenStudy (kendricklamar2014):

5*6=30 18*17=306

OpenStudy (anonymous):

I think it's B. The probabilities change every time a bill is pulled out. The probability of pulling out a 1 first is 5 out of 18 which is .227 then the probability of pulling out a 10 is 6 out of 17 because you don't replace the bill. since the events are independent of each other, you get .35, then you multiply them to get .098 which in fraction form is answer B.

OpenStudy (kropot72):

\[\frac{30}{306}=\frac{5}{51}\]

OpenStudy (kendricklamar2014):

Thanks @kropot72

OpenStudy (kropot72):

Dividing the numerator an the denominator of the left hand fraction by 6.

OpenStudy (kropot72):

You're welcome :)